How would solve: n exponent2-6n=-8?
Answer by ReadingBoosters(3246)

Starting with n^2 - 6n = - 8,
Set the equation equal to 0
n^2 - 6n + 8 = -8 + 8
n^2 - 6n + 8 = 0
Solve the quadratic by finding multiples of 8 which add to -6
4 and 2 work: 4*2=8
Both must be negative to get -6, where -4 + - 2 = -6
It works because -4*-2=8
So,
(n - 4)(n - 2)
Therefore,
n-4=0
n - 4 + 4 = 0 + 4
n = 4
n - 2 = 0
n - 2 + 2 = 0 + 2
n = 2
Ultimately,
n = 4 and n = 2